OpenStudy (anonymous):

How do I solve 2y'' - 3y' - 2y = 13 - 2x^2? Please, its a general solution.

OpenStudy (anonymous):

Isn't that 13 - 4 on the right side?

OpenStudy (anonymous):

2x^2 pls

sam (.sam.):

?

sam (.sam.):

find general solution?

OpenStudy (anonymous):

Yes Sam

sam (.sam.):

try let \[y\text{ = }e^{k x}\]

OpenStudy (anonymous):

OK Sam, am very new to this n hope you help with the answer

sam (.sam.):

\[2 y''-3 y'-2 y=13-2^2\] \[2 \frac{d^2\text{}}{dx^2}\left(e^{k x}\right)-3 \frac{d\text{}}{dx}\left(e^{k x}\right)-2 e^{k x}\text{ = }0\] sub \[\text{}\frac{d^2\text{}}{dx^2}\left(e^{k x}\right)\text{ = }k ^2 e^{k x}\text{ and }\frac{d\text{}}{dx}\left(e^{k x}\right)\text{ = }k e^{k x}\] \[2k ^2 e^{k x}-3 k e^{k x}-2 e^{k x}\text{ = }0\] factor \[\left(2 k^2-3 k -2\right) e^{k x}\text{ = }0\] \[2 k ^2-3k -2\text{ = }0\] \[k=-\frac{1}{2}\text{ or }k =2\]

OpenStudy (anonymous):

Sam how do I edit my question?

sam (.sam.):

there's a button beside your question

OpenStudy (anonymous):

= 13 - 2x^2. 2 ex raised to power 2 pls

OpenStudy (anonymous):

Thanx Sam, I ve done that. Pls I have an exam on that and even IVPs

OpenStudy (anonymous):

= 13 - 2x^2. 2 ex raised to power 2 pls

OpenStudy (anonymous):

Sam how do I edit my question?

OpenStudy (anonymous):

OK Sam, am very new to this n hope you help with the answer

OpenStudy (anonymous):

find out the complimentary function and the particular integral...... add the two....... y=C.F+P.I

OpenStudy (anonymous):

Okay Thanks

sam (.sam.):

did you tried wolfram?

OpenStudy (anonymous):

no i havent its been tough,,